Fermi National Laboratory


Accelerator Update

Monday August 16
The day shift began with the Tevatron (TeV) turn on progressing, with the Antiproton Source (Pbar) stacking, and with most of the other machines conducting studies.

At 9:06 AM, a Booster power supply (B:MV2) tripped off for the first of three times in twenty-five minutes. At 9:38 AM, technicians reported that there was a problem with the bottom power supply. Operations will keep resetting the supply until the TeV completes its load of protons, then techs will replace the supply.

At 10:19 AM, the Main Injector (MI) RF system tripped off. It reset okay.

At 11:16 AM, Operations reported that a Booster RF station (BRF5) tripped off for the first of four times in twenty minutes. At 11:42 AM, Operations bypassed the station. RF techs had the station repaired and back on line fifteen minutes later.

At 11:45 AM, Operations reported a Booster beam problem due to a feed forward system. Investigations pointed to a power supply regulation problem with GMPS. Experts will continue the investigation.

Operations established a 4x0 study store in the TeV at 11:56 AM.

Operations put the TeV into shot setup at 12:36 PM.

At 1:10 PM, Operations noticed an alarm that stated a MI flying wire (MW702) was in the beam. An expert investigated and confirmed that it was in the beam. Operators moved it to the out position and put it in the proper mode to be discovered by the MI sequencer.

At 2:04 PM, Operations reported that BRF5 was still sparking and tripping.

As if shot setup wasn't being troubled enough by everything else, at 2:25 PM, GMPS stopped working. The techs investigating the regulation problem wiggled a wire.

Booster power supply techs took advantage of the GMPS problem at 2:32 PM to replace the B:MV2 power supply, and the RF techs worked on BRF5.

Operations reported that the Booster BRF5 was back online at 2:53 PM.

At 2:56 PM, Operations reported continuing trouble with the TeV shot setup: terrible emittances, poor coalescing, and with little beam actually making it into the TeV. Investigations showed a phase lock problem between the Accumulator and MI and a ramp problem.

Operations tried a 2mA transfer of antiprotons at 3:47 PM to see if everything was working properly. Operators terminated the beam in the TeV at 3:51 PM and then put the TeV through a dry squeeze at 3:57 PM.

Operations established store 3736 at 6:13 PM with an initial luminosity of 60.19E30.

Pbar resumed stacking at 6:20 PM.

Operations resumed sending beam to the Meson MTest and MCenter experiments at 6:21 PM.

At 7:58 PM, BRF13 tripped off for the first of two times in six minutes. It reset okay.

A Pbar expert began a study period at 8:15 PM.

An Instrumentation tech began working at 8:22 PM on the MiniBooNE BPM system. He replaced a timing chassis and a 5-volt power supply.

The D-Zero (D0) experimenters reported at 9:49 PM that they were inserting the Roman pots. At 10:46 PM, Operations reported that they noticed a beam problem around the same time as the pot insertion, but the two didn't correlate directly. They also saw a movement on the motion detector as a big truck passed, but again, though close, it didn't correlate exactly.

Pbar studies ended at 11:35 PM. Pbar resumed normal stacking.

Tuesday August 17
The midnight shift began with Operations monitoring store 3736, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

At 4:58 AM, the interlocks for Meson MC8 enclosure tripped. Operators search and secured the area and had beam back to the experiment by 5:19 AM.

A MI expert began a slipstacking study at 8:43 AM. A Pbar expert began a Debuncher/Accumulator study at the same time.

At 8:56 AM, Operations reported a TeV beam problem. A TeV expert investigated, but could find no problem. The problem persisted for about two hours and then beam operations returned to normal.

MI slipstacking studies ended at 9:58 AM. They then began NuMI studies.

Pbar ended its study around 10:06 AM.

At 10:21 AM, MiniBooNE began tripping off due to beam problems. Operators tuned the beam line, but this didn't help. At 10:55 AM, experts got involved.

Pbar studies finished at 12:10 PM.

MI NuMI studies ended at 12:46 PM.

Operations reported at 1:10 PM that slipstacking for MI and Pbar would be the normal mode of operations from now on.

At 1:37 PM, Booster stopped accelerating beam. The reason was due to GMPS, but there was no bad status. Experts investigated and had beam back by 2:10 PM, but then it was gone ten minutes later. It came back on at 2:53 PM.

Around 2:06 PM, a Linac RF station (KRF3) tripped off. It reset okay, but then it tripped off again at 2:52 PM along with KRF6 & 7. They reset okay.

At 3:01 PM, after the Linac RF problems were resolved, Operations had trouble resuming stacking due to a Debuncher kicker (D:IKIK) toggling between remote and local operations. An expert investigated but said he couldn't do anything while Operations was using the kicker.

A Pbar expert began another stacking study at 4:04 PM.

TeV experts began an end-of-store study at 5:07 PM.

The Pbar study ended at 5:17 PM.

Around 5:42 PM, Operations terminated store 3736 and prepared the D0 collision hall for access.

At 6:36 PM, MIRF17 tripped off twice in two minutes.

At 6:46 PM, MIRF17 tripped off for the first of many times over the next hour.

At 6:50 PM, Operations reported that the heavy rain was finding its way into the end of Linac where the MuCool RF penetrations enter the tunnel. Operators took care of the water.

D0 experimenters came out of access at 7:17 PM.

Operations put the TeV into shot setup at 8:25 PM. Switchyard (SY) beam gets disabled during shot setup.

At 8:52 PM, Operations reported a problem with a part of the shot setup sequencer. Operators contacted an expert for help.

At 9:03 PM, MI had some vacuum alarms. Everything came back to normal after a few minutes and the TeV resumed shot setup.

At 9:41 PM, the TeV lost its ramp. Operators contacted an expert who came in and replaced a TeV sector B2 power supply by 10:58 PM.

Operations restarted the shot setup at 10:47 PM.

At 11:06 PM, the TeV suffered a sector B2 quench possibly due to the B2 power supply replacement.

Wednesday August 18
The midnight shift began with Pbar stacking, and with MiniBooNE, MTest, and MCenter taking beam.

Operations put the TeV through a dry squeeze at 12:46 AM.

At 1:11 AM, the cryo system techs at CHL reported a stuck cryo valve in sector A1. This valve regulated the Dewar level for the A1 system; this control was passed to another device.

Operations established store 3739 at 3:42 AM with an initial luminosity of 76.45E30.

Pbar resumed stacking.

At 7:13 AM, a Booster vacuum burst tripped off BRF7 & 8. The both reset okay.

Pbar resumed stacking at 7:19 AM.

Plans for Wednesday
The plans for today are to continue with stacking and storing. Pbar experts will be allowed to conduct a reverse protons study this afternoon. The plan for tomorrow is to end the store in the morning and allow D0 to make an hour access.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/18/2004   email Fermilab